OpenStudy (anonymous):

solve the system x+y=12 x-y=2

OpenStudy (amistre64):

7 and 5

OpenStudy (anonymous):

grrrr

OpenStudy (amistre64):

lol....

OpenStudy (anonymous):

my fingers are just not that fast

OpenStudy (anonymous):

x+y=12--(1) x-y=2----(2) (1)+(2)-- 2x=14 x=7,y=5

OpenStudy (amistre64):

it had to be 2 numbers that made 12 and were 2 apart; 8 and 4? nope... 7 and 5? yep lol

OpenStudy (anonymous):

yeah visual